DumbChick:

Jorge and Mary are both filling cups with fruit punch in preparation for a party. After 3 minutes, Jorge had 53 cups left to be filled, and after 5 minutes, he had 25 cups left to be filled. Mary’s progress is shown in the table of values below. Assuming that both Jorge and Mary are filling cups at a constant rate, which statement is correct? Mary’s Cup-Filling Progress Time (minutes) 0.5 1 1.5 2 2.5 Cups Remaining 99 81 63 45 27 Jorge is filling 14 cups per minute, which is faster than Mary. Jorge is filling 28 cups per minute, which is faster than Mary. Mary is filling 18 cups per minute, which is faster than Jorge. Mary is filling 36 cups per minute, which is faster than Jorge.

jhonyy9:

so look please - given that Jorge after 3 minutes had 53 cups left to be filled and after 5 minutes ,he had 25 cups left to be filled so from this you can calcule that in these 5 - 3 = 2 minutes Jorge filled 53 -25 = 28 cups from what result that he filled 28/2 = 14 cups/minute - so in this above wrote way please try calcule Mary how many cups left to be filled in one minute hope helped understandably easy

jhonyy9:

@supie opinion pls - ty